Prove that the function f:R2 to R defined by f(x,y) =

{ xy/(x2+y2) if (x,y) not equal to (0,0)

0, if (x,y) = (0,0)

Is not continuous at (0,0)


1
Expert's answer
2021-03-17T09:47:10-0400

To verify whether this function is continuous at "(0;0)" we need to study the limit of "\frac{xy}{x^2+y^2}" when "(x,y)\to(0,0)". When calculating limits of this type (especially in "\mathbb{R}^2") , it is very useful to pass into the polar coordinate system, where the limit "(x,y)\to (0,0)" is replaced by a limit "r\to 0" independently of "\theta". In polar coordinates "x=r\cos \theta, y=r\sin \theta" :

"\lim_{x,y\to 0} \frac{xy}{x^2+y^2}=\lim_{r\to 0} \frac{r\cos{\theta}\text{ } r\sin{\theta}}{r^2}=\lim_{r\to 0}\sin\theta \cos\theta"

And we see that this limit depends on "\theta" and thus the limit "r\to 0" does not exist, so the function "f(x,y)" is not continuous at "(0,0)". For more visibility, we can calculate the limit along "x=0, y\neq 0" which gives "f(0,y)=0\to 0" and the limit along "x=y\neq 0" which gives "f(t,t)=t^2/(2t^2)=1/2\to 1/2\neq 0".